# Canonical bases of tensor products of integrable highest weight modules arising from framed constructions

### v2

@@ 4.1. Framed construction of $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$
Let $\tilde{\mathbf{B}}$ be the canonical basis of $\tilde{\mathbf{f}}$, see 2.5.1.
@@
For any $b\in\tilde{\mathbf{B}}$ and $x\in\tilde{\mathbf{f}}$, suppose that $x=\sum_{b^{\prime}\in\tilde{\mathbf{B}}}c_{b^{\prime}}b^{\prime}$, then we say $b$ appears in $x$ if $c_{b}\not=0$. We define the subset $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\subset\tilde{\mathbf{B}}$ consisting of basis elements which appear in $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ for any $\boldsymbol{\nu}_{1}\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}\in\mathcal{V}_{\nu_{2}}$ and $\nu_{1},\nu_{2}\in\mathbb{N}[I]$.
@@
where $c_{b^{\prime}},d_{b^{\prime\prime}}\in\mathbb{N}[v,v^{-1}]$. If $b\in\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, suppose that $b$ appears in $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ for some $\boldsymbol{\nu}_{1}\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}\in\mathcal{V}_{\nu_{2}}$ and $\nu_{1},\nu_{2}\in\mathbb{N}[I]$. Then $\pi_{i,n}(b)$ and those $b^{\prime}$ appear in $\theta_{(ni,\boldsymbol{\nu}_{1})}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$, where $(ni,\boldsymbol{\nu}_{1})\in\mathcal{V}_{ni+\nu_{1}}$ is the sequence obtained by adding $ni$ before the sequence $\boldsymbol{\nu}_{1}$, and so we have $\pi_{i,n}(b)\in\tilde{\mathbf{B}}_{i,n}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $b^{\prime}\in\tilde{\mathbf{B}}_{i,\geqslant n+1}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Conversely, if $b\in\tilde{\mathbf{B}}_{i,0}$ and $\pi_{i,n}(b)\in\tilde{\mathbf{B}}_{i,n}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, suppose that $\pi_{i,n}(b)$ appears in $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ for some $\boldsymbol{\nu}_{1}=(a_{1}i_{1},...,a_{n}i_{n})\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}=(a^{\prime}_{1}j_{1},...,a^{\prime}_{m}j_{m})\in\mathcal{V}_{\nu_{2}}$ and $\nu_{1},\nu_{2}\in\mathbb{N}[I]$. We identify $\theta_{\lambda}$ with $\theta_{(\langle i,\lambda\rangle i^{\prime})_{i\in I}}$ for $(\langle i,\lambda\rangle)i^{\prime})_{i\in I}\in\mathcal{V}_{|\theta_{\lambda}|}$ in arbitrary order. By [13, Section 9.2.11], we have
@@
where the sum is taken over $\boldsymbol{\nu}\in\mathcal{V}_{ni},\boldsymbol{\nu}^{\prime}\in\mathcal{V}_{\nu_{1}+|\theta_{\lambda}|+\nu_{2}-ni}$ such that
$\boldsymbol{\nu}+\boldsymbol{\nu}^{\prime}=(a_{1}i_{1},...,a_{n}i_{n},(\langle i,\lambda\rangle)i^{\prime})_{i\in I},a^{\prime}_{1}j_{1},...,a^{\prime}_{m}j_{m}),$
and $d(\boldsymbol{\nu},\boldsymbol{\nu}^{\prime})\in\mathbb{Z}$. Notice that such $\boldsymbol{\nu}^{\prime}$ must have the form $(\boldsymbol{\nu}^{\prime}_{1},(\langle i,\lambda\rangle)i^{\prime})_{i\in I},\boldsymbol{\nu}^{\prime}_{2})$, where $\boldsymbol{\nu}^{\prime}_{1}\in\mathcal{V}_{\nu^{\prime}_{1}},\boldsymbol{\nu}^{\prime}_{2}\in\mathcal{V}_{\nu^{\prime}_{2}}$ and $\nu^{\prime}_{1},\nu^{\prime}_{2}\in\mathbb{N}[I]$ such that $\nu^{\prime}_{1}+\nu^{\prime}_{2}=\nu_{1}+\nu_{2}-ni$, thus $\theta_{\boldsymbol{\nu}^{\prime}}=\theta_{\boldsymbol{\nu}^{\prime}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}^{\prime}_{2}}$. Notice that $\theta_{\boldsymbol{\nu}}=c_{\boldsymbol{\nu}}\theta_{i}^{(n)}$ for some $c_{\boldsymbol{\nu}}\in\mathbb{N}[v,v^{-1}]$, so $b$ and those $b^{\prime\prime}$ appear in $\sum c_{\boldsymbol{\nu}}\theta_{\boldsymbol{\nu}^{\prime}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}^{\prime}_{2}}$, and moreover, they appear in some $\theta_{\boldsymbol{\nu}^{\prime}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}^{\prime}_{2}}$, and so we have $b\in\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant 1}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Hence $\pi_{i,n}$ can be restricted to be the bijection from $\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ to $\tilde{\mathbf{B}}_{i,n}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. The other statements can be proved dually. ∎
@@
We make an induction on $\mathrm{tr}\,\nu_{1}+\mathrm{tr}\,\nu_{2}\in\mathbb{N}$ and descending inductions on
@@
$\bullet$ If $\mathrm{tr}\,\nu_{1}+\mathrm{tr}\,\nu_{2}=0$, then $b$ appears in $\theta_{\lambda}$, and we have $b=\theta_{\lambda}\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
By Theorem 2.6, $\bigcup_{i\in I}\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}$ is a basis of $\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}$, and $\bigcup_{i\in I}(\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}\cup\tilde{\mathbf{B}}_{i^{\prime},\geqslant\langle i,\lambda\rangle+1}^{\sigma})$ is a basis of $\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}+\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i^{\prime}}^{\langle i,\lambda\rangle+1}$. Moreover, by Proposition 4.3, we know $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\cap\bigcup_{i\in I}\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}$ is a basis of $\mathbf{f}\theta_{\lambda}\mathbf{f}\cap\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}$, and $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\cap\bigcup_{i\in I}(\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}\cup\tilde{\mathbf{B}}_{i^{\prime},\geqslant\langle i,\lambda\rangle+1}^{\sigma})$ is a basis of $\mathbf{f}\theta_{\lambda}\mathbf{f}\cap(\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}+\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i^{\prime}}^{\langle i,\lambda\rangle+1})$. Notice that $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\cap\bigcup_{i\in I}\tilde{\mathbf{B}}_{i^{\prime},\geqslant\langle i,\lambda\rangle+1}^{\sigma}=\varnothing$, and so
@@
[…]
where $\tau$ is the isomorphism swapping two factors, and $\pi_{\xi\odot\lambda}$ is the natural projection. It is clear that $\chi^{\prime}$ is surjective, and $\chi^{\prime}((\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\xi\rangle+1})\otimes\mathbf{f}\theta_{\lambda}\mathbf{f})=0$, so $\chi^{\prime}$ induced a well-defined surjective linear map $\chi:\Lambda_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f}\rightarrow\Lambda_{\xi,\lambda}$, that is, $\chi(\Lambda_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f})=\Lambda_{\xi,\lambda}$. Let $S$ be the subspace of $\mathbf{f}\theta_{\lambda}\mathbf{f}$ spanned by $\tilde{\mathbf{B}}(0,\lambda)$. We claim that
@@
It is clear that $\chi(\Lambda_{\xi}\otimes S)\subset\chi(\Lambda_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f})$. For any $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\setminus\tilde{\mathbf{B}}(0,\lambda)$, we have $b\in\bigcup_{i\in I}\tilde{\mathbf{B}}_{i,\geqslant 1}^{\sigma}$. We make an induction on $\mathrm{tr}\,|b|$ and a descending induction on $t_{i}^{\sigma}(b)\in[1,|b|_{i}]$ for any $i\in I$ to prove that $\chi(m\otimes b)\in\chi(\Lambda_{\xi}\otimes S)$ for any $m\in\Lambda_{\xi}$. Suppose that $b\in\tilde{\mathbf{B}}_{i,\geqslant 1}^{\sigma}$, and so $t_{i}^{\sigma}(b)\geqslant 1$. By Theorem 2.6 and Lemma 4.2, there exists a unique $b^{\prime}\in\tilde{\mathbf{B}}_{i,0}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that
@@
for any $\xi^{\prime},\lambda^{\prime}\in X$. Hence any surjective linear maps from $\Lambda_{\xi,\lambda}$ to $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ preserving weight subspaces are linear isomorphism. ∎
@@
We remark that the isomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ does not preserve the inner products $(,)_{\xi\odot\lambda}$ and $(,)_{\xi,\lambda}$, see Lemma 3.10. As a result, we can not use the almost orthogonal property of the canonical basis $\tilde{\mathbf{B}}$ to prove Lemma 4.6 directly.
@@
We use Lemma 2.4 and Theorem 2.9 as follows. Recall that $\mathcal{L}_{\xi,\lambda}$ is the $\mathbb{Z}[v^{-1}]$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$, and so $\mathcal{A}\otimes_{\mathbb{Z}[v^{-1}]}\mathcal{L}_{\xi,\lambda}$ is the $\mathcal{A}$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$.
For any $b\in\tilde{\mathbf{B}}(\xi,\lambda)\subset\tilde{\mathbf{B}}$, we have $b\in{{}_{\mathcal{A}}\tilde{\mathbf{f}}}$, see Theorem 2.5. Since the comultiplication of $\tilde{\mathbf{f}}$ maps ${{}_{\mathcal{A}}\tilde{\mathbf{f}}}$ to ${{}_{\mathcal{A}}\tilde{\mathbf{f}}}\otimes_{\mathcal{A}}{{}_{\mathcal{A}}\tilde{\mathbf{f}}}$, see [13, Proposition 14.2.6(a)], and the right multiplication by $\theta_{\lambda}$ maps ${{}_{\mathcal{A}}\mathbf{f}}$ to ${{}_{\mathcal{A}}\tilde{\mathbf{f}}}$, by the definitions of $\varphi$ and $\psi$, we have $\varphi\pi_{\xi\odot\lambda}(b)\in\mathcal{A}\otimes_{\mathbb{Z}[v^{-1}]}\mathcal{L}_{\xi,\lambda}$. By Lemma 4.6, we have $(\varphi\pi_{\xi\odot\lambda}(b),\varphi\pi_{\xi\odot\lambda}(b))_{\xi,\lambda}\in 1+v^{-1}\mathbf{A}$. By Lemma 2.4, there exists $b_{1}\otimes b_{2}\in\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$ such that
$\varphi\pi_{\xi\odot\lambda}(b)=\pm b_{1}\otimes b_{2}\ \mathrm{mod}\,v^{-1}L(\Lambda_{\xi}\otimes\Lambda_{\lambda}),$
where $L(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ is the $\mathbf{A}$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$. Since $\varphi\pi_{\xi\odot\lambda}(b)\in\mathcal{A}\otimes_{\mathbb{Z}[v^{-1}]}\mathcal{L}_{\xi,\lambda}$ and $\mathcal{A}\cap\mathbf{A}=\mathbb{Z}[v^{-1}]$, we have
$\varphi\pi_{\xi\odot\lambda}(b)=\pm b_{1}\otimes b_{2}\ \mathrm{mod}\,v^{-1}\mathcal{L}_{\xi,\lambda}.$

### v3

@@ 4.1. Framed construction of $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$
Let $\tilde{\mathbf{B}}$ be the canonical basis of $\tilde{\mathbf{f}}$ defined in [13, Chapter 14].
@@
For any $b\in\tilde{\mathbf{B}}$ and $x\in\tilde{\mathbf{f}}$, suppose that $x=\sum_{b^{\prime}\in\tilde{\mathbf{B}}}c_{b^{\prime}}b^{\prime}$. Then we say $b$ appears in $x$ if $c_{b}\not=0$. We define the subset $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\subset\tilde{\mathbf{B}}$ consisting of $b\in\tilde{\mathbf{B}}$ which appear in $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ for some $\boldsymbol{\nu}_{1}\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}\in\mathcal{V}_{\nu_{2}}$ and $\nu_{1},\nu_{2}\in\mathbb{N}[I]$.
@@
where $c_{b^{\prime}},d_{b^{\prime\prime}}\in\mathbb{N}[v,v^{-1}]$. If $b\in\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, suppose that $b$ appears in $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ for some $\boldsymbol{\nu}_{1}\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}\in\mathcal{V}_{\nu_{2}}$ and $\nu_{1},\nu_{2}\in\mathbb{N}[I]$, that is, if we write $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}=\sum_{b_{0}\in\tilde{\mathbf{B}}}k_{b_{0}}b_{0}$, then $k_{b}\not=0$. Thus we have
$\theta_{i}^{(n)}\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}=\theta_{i}^{(n)}(k_{b}b+\sum_{b_{0}\in\tilde{\mathbf{B}},b_{0}\not=b}k_{b_{0}}b_{0})=k_{b}\pi_{i,n}(b)+\sum_{b^{\prime}\in\tilde{\mathbf{B}}_{i,\geqslant n+1}}k_{b}c_{b^{\prime}}b^{\prime}+\sum_{b_{0}\in\tilde{\mathbf{B}},b_{0}\not=b}k_{b_{0}}\theta_{i}^{(n)}b_{0}.$
Since the monomial $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ is a product of divided powers which are canonical basis elements, by the positivity of $\tilde{\mathbf{B}}$ with respect to the multiplication, we have $k_{b},k_{b_{0}}\in\mathbb{N}[v,v^{-1}]$, and $\theta_{i}^{(n)}b_{0}$ is a $\mathbb{N}[v,v^{-1}]$-linear combination of canonical basis elements in $\tilde{\mathbf{B}}$. Hence $\pi_{i,n}(b)$ and those $b^{\prime}\in\tilde{\mathbf{B}}_{i,\geqslant n+1}$ such that $c_{b^{\prime}}\not=0$, appear in $\theta_{(ni,\boldsymbol{\nu}_{1})}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$, where $(ni,\boldsymbol{\nu}_{1})\in\mathcal{V}_{ni+\nu_{1}}$ is the sequence obtained by adding $ni$ before the sequence $\boldsymbol{\nu}_{1}$, and so we have $\pi_{i,n}(b)\in\tilde{\mathbf{B}}_{i,n}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $b^{\prime}\in\tilde{\mathbf{B}}_{i,\geqslant n+1}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Conversely, if $b\in\tilde{\mathbf{B}}_{i,0}$ and $\pi_{i,n}(b)\in\tilde{\mathbf{B}}_{i,n}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, suppose that $\pi_{i,n}(b)$ appears in $\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}$ for some $\boldsymbol{\nu}_{1}=(a_{1}i_{1},...,a_{n}i_{n})\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}=(a^{\prime}_{1}j_{1},...,a^{\prime}_{m}j_{m})\in\mathcal{V}_{\nu_{2}}$ and $\nu_{1},\nu_{2}\in\mathbb{N}[I]$. We identify $\theta_{\lambda}$ with $\theta_{(\langle i,\lambda\rangle i^{\prime})_{i\in I}}$ for $(\langle i,\lambda\rangle)i^{\prime})_{i\in I}\in\mathcal{V}_{|\theta_{\lambda}|}$ in arbitrary order. By [13, Section 9.2.11], we have
@@
where the sum is taken over $\boldsymbol{\nu}=(x_{1},...,x_{n+m+|I|})\in\mathcal{V}_{ni},\boldsymbol{\nu}^{\prime}=(y_{1},...,y_{n+m+|I|})\in\mathcal{V}_{\nu_{1}+|\theta_{\lambda}|+\nu_{2}-ni}$ such that
$(x_{1}+y_{1},...,x_{n+m+|I|}+y_{n+m+|I|})=(a_{1}i_{1},...,a_{n}i_{n},(\langle i,\lambda\rangle)i^{\prime})_{i\in I},a^{\prime}_{1}j_{1},...,a^{\prime}_{m}j_{m}),$
and $d(\boldsymbol{\nu},\boldsymbol{\nu}^{\prime})\in\mathbb{Z}$. Notice that such $\boldsymbol{\nu}^{\prime}$ must have the form $(\boldsymbol{\nu}^{\prime}_{1},(\langle i,\lambda\rangle)i^{\prime})_{i\in I},\boldsymbol{\nu}^{\prime}_{2})$, where $\boldsymbol{\nu}^{\prime}_{1}\in\mathcal{V}_{\nu^{\prime}_{1}},\boldsymbol{\nu}^{\prime}_{2}\in\mathcal{V}_{\nu^{\prime}_{2}}$ and $\nu^{\prime}_{1},\nu^{\prime}_{2}\in\mathbb{N}[I]$ such that $\nu^{\prime}_{1}+\nu^{\prime}_{2}=\nu_{1}+\nu_{2}-ni$, thus $\theta_{\boldsymbol{\nu}^{\prime}}=\theta_{\boldsymbol{\nu}^{\prime}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}^{\prime}_{2}}$. Notice that $\theta_{\boldsymbol{\nu}}=c_{\boldsymbol{\nu}}\theta_{i}^{(n)}$ for some $c_{\boldsymbol{\nu}}\in\mathbb{N}[v,v^{-1}]$, so $b$ and those $b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant 1}$ such that $d_{b^{\prime\prime}}\not=0$ appear in $\sum c_{\boldsymbol{\nu}}\theta_{\boldsymbol{\nu}^{\prime}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}^{\prime}_{2}}$, and moreover, they appear in some $\theta_{\boldsymbol{\nu}^{\prime}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}^{\prime}_{2}}$, and so we have $b\in\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant 1}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Hence $\pi_{i,n}$ can be restricted to be the bijection from $\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ to $\tilde{\mathbf{B}}_{i,n}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. The other statements can be proved dually. ∎
@@
We use induction on $\mathrm{tr}\,\nu_{1}+\mathrm{tr}\,\nu_{2}\in\mathbb{N}$ and descending inductions on
@@
$\bullet$ If $\mathrm{tr}\,\nu_{1}+\mathrm{tr}\,\nu_{2}=0$, then $b$ appears in $\theta_{\lambda}$. We claim that $\theta_{\lambda}\in\tilde{\mathbf{B}}_{i^{\prime},0}$. Indeed, it is clear that $1\in\tilde{\mathbf{B}}$. For any $i\in I$, by Theorem 2.6, we have $\theta_{i^{\prime}}^{(\langle i,\lambda\rangle)}=\theta_{i^{\prime}}^{(\langle i,\lambda\rangle)}1=\pi_{i^{\prime},\langle i,\lambda\rangle}(1)\in\tilde{\mathbf{B}}$. Notice that we have $\theta_{i^{\prime}}^{(\langle i,\lambda\rangle)}\in\bigcap_{j\in I,j\not=i}\tilde{\mathbf{B}}_{j^{\prime},0}$, by repeating above process, we obtain $\theta_{\lambda}\in\tilde{\mathbf{B}}$. Hence $b=\theta_{\lambda}\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
By Theorem 2.6, we know that $\bigcup_{i\in I}\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}$ is a basis of $\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}$, and $\bigcup_{i\in I}(\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}\cup\tilde{\mathbf{B}}_{i^{\prime},\geqslant\langle i,\lambda\rangle+1}^{\sigma})$ is a basis of $\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}+\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i^{\prime}}^{\langle i,\lambda\rangle+1}$. Moreover, by Proposition 4.3, we know $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\cap\bigcup_{i\in I}\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}$ is a basis of $\mathbf{f}\theta_{\lambda}\mathbf{f}\cap\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}$, and $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\cap\bigcup_{i\in I}(\tilde{\mathbf{B}}_{i,\geqslant\langle i,\xi\rangle+1}^{\sigma}\cup\tilde{\mathbf{B}}_{i^{\prime},\geqslant\langle i,\lambda\rangle+1}^{\sigma})$ is a basis of $\mathbf{f}\theta_{\lambda}\mathbf{f}\cap(\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1}+\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i^{\prime}}^{\langle i,\lambda\rangle+1})$. Notice that $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\cap\bigcup_{i\in I}\tilde{\mathbf{B}}_{i^{\prime},\geqslant\langle i,\lambda\rangle+1}^{\sigma}=\varnothing$, and so
@@
$\chi^{\prime}:M_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f}=\mathbf{f}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f}\xrightarrow{\tau}\mathbf{f}\theta_{\lambda}\mathbf{f}\otimes\mathbf{f}\xrightarrow{\mathrm{multiplication\ of}\ \tilde{\mathbf{f}}}\mathbf{f}\theta_{\lambda}\mathbf{f}\xrightarrow{\pi_{\xi\odot\lambda}}\Lambda_{\xi,\lambda},$
where $\tau$ is the isomorphism swapping two factors, and $\pi_{\xi\odot\lambda}$ is the natural projection. It is clear that $\chi^{\prime}$ is surjective, and $\chi^{\prime}((\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\xi\rangle+1})\otimes\mathbf{f}\theta_{\lambda}\mathbf{f})=0$, so $\chi^{\prime}$ induces a well-defined surjective linear map $\chi:\Lambda_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f}\rightarrow\Lambda_{\xi,\lambda}$, that is, $\chi(\Lambda_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f})=\Lambda_{\xi,\lambda}$. Let $S$ be the subspace of $\mathbf{f}\theta_{\lambda}\mathbf{f}$ spanned by $\tilde{\mathbf{B}}(0,\lambda)$. We claim that
@@
It is clear that $\chi(\Lambda_{\xi}\otimes S)\subset\chi(\Lambda_{\xi}\otimes\mathbf{f}\theta_{\lambda}\mathbf{f})$. For any $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})\setminus\tilde{\mathbf{B}}(0,\lambda)$, we have $b\in\bigcup_{i\in I}\tilde{\mathbf{B}}_{i,\geqslant 1}^{\sigma}$. We use induction on $\mathrm{tr}\,|b|$ and a descending induction on $t_{i}^{\sigma}(b)\in[1,|b|_{i}]$ for any $i\in I$ to prove that $\chi(m\otimes b)\in\chi(\Lambda_{\xi}\otimes S)$ for any $m\in\Lambda_{\xi}$. Suppose that $b\in\tilde{\mathbf{B}}_{i,\geqslant 1}^{\sigma}$, and so $t_{i}^{\sigma}(b)\geqslant 1$. By Theorem 2.6 and Lemma 4.2, there exists a unique $b^{\prime}\in\tilde{\mathbf{B}}_{i,0}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that
@@
for any $\xi^{\prime},\lambda^{\prime}\in X$. Hence any surjective $\mathbf{U}^{0}$-module homomorphisms from $\Lambda_{\xi,\lambda}$ to $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ are isomorphisms. ∎
@@
We remark that the isomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ does not preserve the bilinear forms $(,)_{\xi\odot\lambda}$ and $(,)_{\xi,\lambda}$, see Lemma 3.10. As a result, we can not use the almost orthonormal property of the canonical basis $\tilde{\mathbf{B}}$ to prove Lemma 4.6 directly.
@@
We use Lemma 2.4 and Theorem 2.9. Recall that $\mathcal{L}_{\xi,\lambda}$ is the $\mathbb{Z}[v^{-1}]$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$, and so $\mathcal{A}\otimes_{\mathbb{Z}[v^{-1}]}\mathcal{L}_{\xi,\lambda}$ is the $\mathcal{A}$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$.